Solving using matrices:
3x+5y+2z=3
-x-y-z=-2
2x-2y+5z=11

Answer by Alan3354(69443)

3x+5y+2z=3
-x-y-z=-2
2x-2y+5z=11
------------
Det:
3(-5-2) -5(-5+2) + 2(2+2) = -21 + 15 + 8
Det = 2
--------
x:
5(-11+10) - 2(-11-4) + 3(-5-2) = -5 + 30 - 21 = 4
x = 4/Det = 4/2
x = 2
------
-y:
3(-11+10) -2(-11+4) +3(-5+2) = -3 +14 -9 = 2
y = -1
------
z:
3(-11-4) -5(-11+4) + 3(2+2) = -45 +35 +12 = 2
z = 1
---------
3x+5y+2z=3 --> 6 - 5 + 2 = 3
-x-y-z=-2 ---> -2 + 1 - 1 = -2
2x-2y+5z=11 --> 4 + 2 + 5 = 11
It checks
Any questions, email me at gsihoutx@aol.com